Circular No. 6350

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 AND GRO J1744–28 K. Borozdin, N. Alexandrovich, and R. Sunyaev, Space Research In-

stitute, Moscow, report on behalf of MIR-KVANT-TTM team: “The TTM experiment onboard the Mir-Kvant Space Station has detected GRS 1739– 278 (IAUC 6348) as a very bright and soft x-ray source. The 2- to 27-keV flux was equal to 200 ± 50 mCrab on Feb. 28, rising to 500 ± 50 mCrab on Mar. 1. The TTM position (α = 17^{h}42^{m}40^{s}, δ = −27^{o}45′.8, equinox 2000.0; uncertainty ± 1′), is inside the SIGMA/Granat error box. The source was weaker than 100 mCrab during previous TTM sessions of GC observation (Feb. 6–7). The spectrum in the band 2–27 keV can be roughly fitted by a power law with spectral index 2.7 ± 0.1 and hydrogen absorption column density (4.1 ± 0.7) × 10^{22} cm^{−2}. However, the two-component (soft component plus hard power law) model with smaller absorption provides better agreement with data obtained on Mar. 1. Such x-ray spectra are typical for accreting blackhole candidates. GRO J1744–28 was observed simultaneously during galactic field observations. The 2- to 27-keV flux was 1650 ± 70 mCrab on Feb. 7, 515 ± 30 mCrab on Feb. 28, and 1220 ± 50 mCrab on Mar. 1. The TTM localized the source at α = 17^{h}44^{m}34^{s}, δ = −28^{o}45′.2 (uncertainty ± 1′); this source has a much harder spectrum in the band 2–27 KeV than does GRS 1739–278.”

COMET C/1996 B2 (HYAKUTAKE) C. Lisse, Goddard Space Flight Center; and K. Dennerl and J. Schmitt,

Max-Planck-Institut für Extraterrestrische Physik, Garching, communicate: “Comet C/1996 B2 will be observed with the x-ray satellite ROSAT. The scheduled observing intervals are as follows: Mar. 26.493–26.529 UT, 26.626–26.634, 26.638–26.662, 26.708–26.729, 26.89–26.902, 27.554–27.591, 27.687–27.695, 27.703–27.724, 27.772–27.789, 27.841–27.856, 28.417–28.44. Simultaneous observations at other wavelengths are encouraged.”

Naked-eye m_{1}, coma-diameter, and tail-length estimates: Mar. 21.93 UT, 0.8, 67′, – (K. Černis, Rukainiai, Lithuania); 22.41, 0.9, 1^{o}.5, 31^{o} (R. Keen, Mt. Thorodin, CO); 22.60, 1.0, 90′, 30^{o} (S. J. O’Meara, Volcano, HI); 22.68, 0.3, 70′, 40^{o} (P. Camilleri, Cobram, Vic., Australia); 23.30, 0.5, 2^{o}, 30^{o} (D. W. E. Green, near S. Carver, MA); 23.45, 0.2, –, 45^{o} (C. S. Morris, Whitaker Peak, CA).

1996 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green